OpenStudy (anonymous):

suppose you choose a marble from a bag containing two red marbles 5 white marble & 3 blue marbles you return the first marble to the bag and then choose again find the probability of red and blue

Directrix (directrix):

Practice this type of problem here: http://www.mathscore.com/math/practice/Object%20Picking%20Probability/

Directrix (directrix):

Begin by noting that since the marbles are replaced, each draw does not depend on previous draws and thus the draws are independent

Directrix (directrix):

P(1 red marble on first pick) = 2/10 = 1/5. P(1 blue marble on second pick) = 3/10.

Directrix (directrix):

P( 1 red ball and then 1 blue ball) = 1/5 * 3/10 = ? @candice2k14